OpenStudy (anonymous):

A rectangular raft of thickness 30cm and mass 120kg floats in water.The submerged depth is 16cm.What is the submerged depth when a student of mass 60kg gets on the raft? Te answer is 24cm but I have no idea how to solve it.Please explain in details.Thanks

OpenStudy (anonymous):

180g = V*rho*g find V that's how much of the raft is submerged.

OpenStudy (anonymous):

I dont get what you did thereCan you explain

OpenStudy (anonymous):

we are looking for depth

OpenStudy (anonymous):

sure, they tell you 120g = V*rho*g they also say h(submerged) = .16 so area = 120/(.16rho)

OpenStudy (anonymous):

use that in the equation: 180g = V*rho*g where V = 120h/(.16rho)

OpenStudy (anonymous):

area = 120/(.16rho) ?

OpenStudy (anonymous):

yes.

OpenStudy (anonymous):

I dont get where area cme from

OpenStudy (anonymous):



OpenStudy (anonymous):

what happend to g?

OpenStudy (anonymous):

should it be v=120h/(0.16rhpg)

OpenStudy (anonymous):

go for it.

OpenStudy (anonymous):

did you get 24 cm?

OpenStudy (anonymous):

wait

OpenStudy (anonymous):

180g = (120h/(.16rho) )* rho*g h= .16*(180/120)

OpenStudy (anonymous):

I am totally confused

OpenStudy (anonymous):

let me go step by step

OpenStudy (anonymous):

(120)g=pvg

OpenStudy (anonymous):

120=pv

OpenStudy (anonymous):

120=pAH

OpenStudy (anonymous):

A=120/(ph)

OpenStudy (anonymous):

good so far.

OpenStudy (anonymous):

180=ph

OpenStudy (anonymous):

pv

OpenStudy (anonymous):

180=pv

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

180=pAH

OpenStudy (anonymous):

180=A=120/(ph).P.h

OpenStudy (anonymous):

180=A.120/(ph).P.h

OpenStudy (anonymous):

i THINK THERE SHOULD BE AN EASIER WAY

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

it is worth 0.5 point

OpenStudy (anonymous):

I already did it the easy way.

OpenStudy (anonymous):

look at it this way if you like mass of the floating object is proportional to volume of water displaced if the volume is something simple, like a rectangular prism, then the height submerged is proportional to volume submerged. m2/m1 = V2/V1 = h2/h1 since the constants will cancel in every case. (edited a few times for clarity)